the length of a rectangle is twice the width. the area is 392 yd^2. find the length and the width
Answer by checkley77(12569)

L=2W
A=LW
392=2W*W
2W^2=392
W^2=392/2
W^2=196
W=SQRT196
W=14 ANS. FOR THE WIDTH.
L=2*14=28 ANS. FOR THE LENGTH.
PROOF:
392=28*14
392=292